OpenStudy (math2400):

Chromium crystallizes with a body-centered cubic unit cell. The radius of a chromium atom is 125 pm. Calculate the density of solid crystalline chromium in g/cm3.

OpenStudy (math2400):

r=125 Cr: 51.9961g/mol * (1/6.022*10^23) = 8.63*10^-23 g/atom 2 atoms per unit cell so multiply and get: 1.762*10^-14 then length=4r/3^1/2 length of body-centered cubic unit cell then cube that (r is 125*10^-12) anything wrong so far??